OpenStudy (anonymous):

Let f(x) = (x^2 -5x - 14)/(x-7). Identify another function that agrees with f(x) at all but one point.

OpenStudy (anonymous):

Could you define a piecewise function? e.g \[f(x)=(x^2-5x-14)/(x-7) for ( -\infty,2) and (2,\infty)\] so the function does not include x=2, yet agrees everywhere else?

OpenStudy (anonymous):

yes, that sounds correct

OpenStudy (anonymous):

wait, shouldn't it be 7 instead of 2? Since it would make the function undefined?

OpenStudy (anonymous):

The original function itself does not contain 7 in domain..So, your answer would be something like this (-infinity,2)U(2,infinity)-{7}

OpenStudy (anonymous):

The domain of the function in the question is R-{7} So, your answer satisfies every value except at x=2

OpenStudy (anonymous):

the function you are looking for is \(g(x)=x-2\)

OpenStudy (anonymous):

scratch that, it is \(g(x)=x+2\) sorry

OpenStudy (anonymous):

Toatally forgot it @satellite73

OpenStudy (anonymous):

*totally